The following information pertains to the most recent quarter at Precious Production Limited. Purchases of raw materials Raw1. Prepare a schedule of cost of goods manufactured. PRECIOUS PRODUCTION Schedule of Cost of Goods Manufactured For the quart2. Prepare an income statement. PRECIOUS PRODUCTION LIMITED Income Statement For the quarter ended xxxx Sales $ 1,896,000 Cos3. Assume that the company produced the equivalent of 14,000 units of product during the year. What was the average cost per6. Assuming the company produced 21,000 fully and partially finished units during the year, determine the cost components of

Please help answer question #3, #4 & #6
Answer #1

3.

Average cost
Direct labor $             17.34 per unit
Insurance $                1.46 per unit

Direct labor = $242800/14000 = $17.34

Insurance = $20400/14000 = $1.46

4.

Average cost per unit Total cost
Direct materials $             24.11 $     385,829
Insurance $                1.28 $         20,400

Direct materials total cost = $337600 x 16000/14000 = $385829

Average cost = $337600/14000 = $24.11

Insurance = $20400/16000 = $1.28 per unit

6.

Direct materials $           67,520
Direct labor $           48,560
Manufacturing overhead $       139,368
Total $       255,448

Direct materials = $337600 x 4200/21000 = $67520

Direct labor = $242800 x 4200/21000 = $48560

Manufacturing overhead = $696840 x 4200/21000 = $139368